Suppose you take a sample of size 4 from a population that is normally distributed with a known standard deviation of 10. You get a sample mean of 20. What is the probability the population mean is 30 or more?

Question 4 options:

0.025

0.05

0.5

0.95

Either 0 or 1, but we don't know which.

Question 5 (1 point)

You want to estimate the proportion of Canadians who have read fewer than 2 books in the past year. You think the population proportion is about 0.65. What is the minimum sample size you need to be able to use a normal approximation in your statistical analysis?

Question 5 options:

29

15

16

28
Answer #1

Given that

sample size n = 4

sample mean = 20.

standard deviation =10

probability the population mean is 30 or more

P( X > 30) = 30 - 20 /10 / √4 = 10 / 5 = 2

Using standard normal table

P ( Z>2 )=1−P ( Z<2 )=1−0.9772=0.0228 = 0.023

  

  *** Dear student we answer one question once post remaining separately****